A rectangular garden has 50m length,40m breadth and 2m wide path running around within the garden,then find the area of the path.
Answer by ankor@dixie-net.com(22740)

A rectangular garden has 50m length,40m breadth and 2m wide path running around within the garden,then find the area of the path.
:
The garden without the path dimensions are reduced by 4 m
46 by 36
:
Garden with path - garden without path = path area
(50 * 40) - (46 * 36) =
2000 - 1656 = 344 sq/m is the area of the path